OpenStudy (anonymous):

Write the equation of the line that is parallel to the line 4x − 3y = −12 and passes through the point (−3, 4).

OpenStudy (anonymous):

anyone who answers right gets medal

OpenStudy (anonymous):

isolate y:\[3y = 4x + 12\] divide both sides by 3:\[y = \frac{ 4 }{ 3 }x + 4\] Now that you have your original equation in the form of y=mx+b, you can see what the slope is, which is 4/3. For your new line to be parallel, it must also have a slope of 4/3. So far you have for your new line: \[y = \frac{ 4 }{ 3 }x + b\] to solve for b, since you know it passes through the point (-3,4), you can substitute those values into your equation: \[4 = \frac{ 4 }{ 3 }(-3) + b\]\[4 = -4 + b\]\[b=8\] So your new line is :\[y = \frac{ 4 }{ 3 }x + 8\]

OpenStudy (anonymous):

thank you